please show full working and check calculations/formulas again before submitting2. (50 pts) In this problem, assume the operational amplifier is ideal and its open-loop gain A=c0. The current sources and v
Answer #1

2L • 2K12 - GV FVV I imp ТСе 5k 1 +5V - VOA 5mA Tokel amit CV 36к. + 2 56lka 3KL IMA -4.5mA سهم 5mA IMA - 3y + tekur 2. m A 32) I Gkr = IMA 5) Vakn = 6kn XIMA = 6 V V XA 3kt 31+ vo va RmA =) V2 -3 -3 V=0 - VL : 6 V y 3) V = 6-6=OV 2ks- 2k1, Ima 2mA 4